Let W be the relationship defined by A W B= 2A + 5B
What is the value of 9W(3 W 1)?
In the problem on the sheet the "W" looks like a pitchfork. I have never seen this symbol before. Please help???

Found 2 solutions by jim_thompson5910, solver91311:
Answer by jim_thompson5910(35256)

Since A W B= 2A + 5B, this means that wherever you see an "A", simply plug in the left value and wherever you see a "B", plug in the right value.


So...

3 W 1 = 2(3) + 5(1) = 6+5 = 11


which means that...

9 W (3 W 1) = 9 W 11 = 2(9) + 5(11) = 18 + 55 =73


So the answer is


9 W (3 W 1) = 73

Answer by solver91311(24713)


The reason that you have never seen the symbol before is that it is made up just for this situation. According to the definition that you provided, it means "multiply the number on the left side of the symbol by 2, then multiply the number on the right of the symbol by 5, and then add the two results. In the example problem you gave, 2 times 3 is 6 and 5 times 1 is 5 and 6 plus 5 is 11, then 2 times 9 is 18 and 5 times 11 is 55 and 18 plus 55 is 73.

I think what your instructor is trying to do is to investigate the Associative and Commutative properties. We know that the standard operations of addition and multiplication are both associative and commutative, but you will find that this new, made-up operation of "fork"-ition is neither commutative or associative in general.

You can prove this to yourself by verifying that



and



John